OpenStudy (anonymous):

For f(x) = 1/x-5 and g(x) = x^2+2, find: A. the expression for g(x). B. substitute the value of g(x) into the function f(x) in place of x to find the value of f(g(x)).

OpenStudy (anonymous):

A? what is it asking for? B... Well its telling you to do a composite function.. SO what we do is this: f(g(x)).. so whats g(x)? g(x) is x^2+2... SO we have f(x^2+2)... So what this means is that for the function of f we plug in x^2+2 for every time we see f... So.... f(g(x))= 1/(x^2+2)-5 Do you know how to simplify that?

OpenStudy (anonymous):

All A says is find the expression for g(x) :O

OpenStudy (anonymous):

is there a way to simplify f(g(x))= 1/(x^2+2)-5

OpenStudy (anonymous):

\[\frac{ 1 }{ x ^{2}+2-5 } = \frac{ 1 }{ x ^{2}-3 }\] then youcan factor that but you get a rational/weird answer.. So i think leaving it like that should be fine

OpenStudy (anonymous):

makes sense! so what would i do about problem A?

OpenStudy (anonymous):

@Gerardo_cast23

OpenStudy (anonymous):

Nothing i guess? Just rewrite g(x)? I googled stuff and cant find anything. Thats a weird/dumb question lol

OpenStudy (anonymous):

okay thanks lol :p